OpenStudy (anonymous):

The total surface area of a right circular cylinder in which te height is twice the radius is 24pi sq. units. find the height.

OpenStudy (anonymous):

SA=2(pi)*(r)*(height)

OpenStudy (campbell_st):

h = 2r then \[24\pi = 2\pi r^2 + 2\pi r \times 2r\] divide by 2 pi \[12 = r^2 + 2r^2\] \[3r^2 =12\] I'll let you finish the problem

OpenStudy (campbell_st):

oops should be \[12 = 5r^2\]